### v4

# A Hybrid Machine Learning Framework for Predicting Hydrogen Storage Capacities: Unsupervised Feature Learning with Deep Neural Networks

Abstract
In this study, we present a sophisticated hybrid machine-learning framework that significantly improves the accuracy of predicting hydrogen storage capacities in metal hydrides. This is a critical challenge due to the scarcity of experimental data and the complexity of high-dimensional feature spaces. Our approach employs the power of unsupervised learning through the use of a state-of-the-art autoencoder. This autoencoder is trained on elemental descriptors obtained from Mendeleev software, enabling the extraction of a meaningful and lower dimensional latent space from the input data. This latent representation serves as the basis for our deep multi-layer perceptron (MLP) model, which consists of five layers and shows good precision in predicting hydrogen storage capacities. Furthermore, our results show very good agreement with the results obtained with density functional theory (DFT). In addition to addressing the limitations caused by limited and unevenly distributed data in the field of hydrogen storage materials, we also focus on discovering new materials that show promising opportunities for hydrogen storage. These materials were identified using both feature-based approaches and predictions generated by a large language model. Finally, our investigation into the effectiveness of transferring weights from the autoencoder to the MLP, in addition to the latent features, suggests that while this strategy slightly improves model performance as indicated by a slightly higher R2 value and lower RMSE, it emphasizes the intricate challenge of adapting pre-trained weights for specific supervised tasks.
keywords:
Hydrogen storage, deep learning
1 Introduction
Climate change, energy security, and the potential depletion of resources propelled by escalating global population and technological advancement constitute pressing challenges for mankind 1, 2. To address these, renewable energy promotion is vital, with hydrogen emerging as the quintessential energy carrier for mobile and stationary applications. Hydrogen not only mitigates environmental harm but also reduces reliance on imported oil for countries devoid of natural resources 3. Various storage techniques exist for hydrogen, including pressurized gas, cryogenic liquid, and solid fuel as a chemical or physical combination with other materials like metal hydrides, complex hydrides, and carbon materials, or produced in-situ in vehicles by methanol reforming 4, 5. Each of these methods possesses unique attributes for hydrogen storage. Storage by absorption as chemical compounds or by adsorption on carbon materials carries safety benefits, necessitating a form of conversion or energy input for hydrogen release. Considerable effort has been invested into new hydrogen-storage systems, including metal, chemical, or complex hydrides, and carbon nanostructures 6.
Hydrogen storage in metal hydrides involves chemically combining hydrogen gas with metal elements, forming metal hydrides. These offer significant safety benefits over gas and liquid storage methods, while also boasting higher hydrogen storage density (6.5 H atoms/cm3 for MgH2) as compared to hydrogen gas (0.99 H atoms/cm3) or liquid hydrogen (4.2 H atoms/cm3) 7. Hence, metal hydride storage is a safe, volume-efficient method suitable for onboard vehicle applications. Hydrogen typically stored within the lattice structure of metal hydrides is attached to metal atoms. Metal-hydrogen combinations yield two types of hydrides, $\alpha$-phase which absorbs some hydrogen, and $\beta$-phase where hydride is fully formed 8. Hydrogen storage in metal hydrides is influenced by various factors and involves several mechanistic steps. The ability of metals to dissociate hydrogen depends on surface structure, morphology, and purity 9. Factors such as the type of metal, type of hydride, and operating conditions impact the hydrogen storage capacity of metal hydrides. Light metals like Li, Be, Na, Mg, B, and Al form numerous metal–hydrogen compounds and are particularly intriguing owing to their lightweight and the large number of hydrogen atoms per metal atom, often in the order of H/M = 2 10, 11, 12, 13, 14.
Many metals interact with hydrogen to form binary hydrides (MHn), but most binary hydrides do not present suitable hydrogen storage properties for serving as a hydrogen carrier. Some intermetallic compounds can also form hydrides with structural formulas of ABxHn, where element A with high hydrogen affinity is usually rare earth or alkaline earth metal forming a stable hydride, while element B has low hydrogen affinity and often forms only unstable hydrides 15, 4. The most extensively studied metal hydrides are AB2 and AB5 for hydrogen storage and fuel cell application 16. AB5-type intermetallic compounds are distinguished by their ease of activation, rapid kinetics of hydrogen absorption and desorption, and relatively high stability of hydrogen sorption properties during cyclic hydrogenation/dehydrogenation 17. Pressure-composition isotherms in H-AB5 systems show a single flat plateau with low H2 absorption-desorption hysteresis, which increases in substituted alloys 18. Ti and Zr are the most common A components in AB2-type compounds, and the B component is usually represented by a transition metal such as Mn, Cr, Fe, or V 17. AB2-type alloys are generally less easily activated than AB5 alloys, and they can be doped with small amounts ($\sim$1 at%) of rare earth elements to facilitate activation 19. They have excellent hydrogenation/dehydrogenation kinetics and cycle stability once activated 20. The AB2 components are less expensive than the AB5 components, but manufacturing of the AB2-type alloys presents some metallurgical challenges due to higher melting temperatures, high component reactivity, and other factors 20, 21. The most studied metal hydrides for hydrogen storage include sodium aluminum hydride, magnesium hydride, aluminum hydride, LaNi5, and ZrV2 22, 23, 18, 24.
The introduction of data-driven materials design 25, 26, 27 has accelerated material discovery, processing, and manufacturing 28, 29, 30. Machine learning (ML) algorithms allow for the creation of new alloys based solely on previously collected data, either publicly available or reported in scientific literature. Metal hydrides are an ideal subject for machine learning algorithm research due to their diverse properties and broad range. Traditional methods for determining the optimized materials class and corresponding metal hydride composition based on desirable properties are difficult, time-consuming, and costly. ML learning techniques, on the other hand, enable rapid, productive, and efficient material class prediction for a specific hydrogen weight percent and operational conditions. Numerous groups have begun to use ML techniques to speed up this screening and gain more physical insight from massive amounts of data 31, 32, 33. Rezakazemi et al. used an adaptive neuro-fuzzy inference system to evaluate the performance of hydrogen-selective mixed matrix membranes under various operational conditions 34. Rahnama et al. predicted the hydrogen storage capacities in various metal hydrides using machine learning algorithms in another study and found that higher temperatures yielded higher hydrogen storage capacities 33, 35. Ahmed et al. recently predicted gravimetric and volumetric hydrogen capacities in metal-organic frameworks (MOFs) using ML algorithms 36. ML methods have been instrumental in predicting various properties of (MOFs) and even designing new structures. This approach has significantly changed the research landscape in this field, allowing for the rapid identification of promising MOF candidates 37, 38 for various applications, including hydrogen storage. In a recent wor, Sun et al. introduced a method that combines meta-learning and high-throughput molecular simulations to predict hydrogen storage in nanoporous materials efficiently. 39
This study proposes a two-stage model for predicting hydrogen storage capacity in metal hydrides. The first stage involves unsupervised learning to extract latent features from the input data, and the second stage employs a multi-layer perceptron (MLP) trained to predict target properties. We developed a deep neural network-based model that can predict the hydrogen storage capacities in metal hydrides. We also identified new hydrogen storage materials using both feature-based approaches as well as using a large language model. The hydrogen storage capacity of these predicted materials was calculated using the above-mentioned approach and further compared with the results from the DFT-based methods. A chemical bonding analysis was performed to comprehend the relationship between storage capacity and the atomic environment, providing insights into the chemical bonds in the materials and their contribution to hydrogen storage capacity.
2 Methods
The architecture of our method is shown in Fig. 1, in a schematic way. We use a hybrid approach that has two parts: in the unsupervised learning part, an autoencoder is trained on the entire dataset, which consists of only unlabeled data, to learn a compact representation of the data in the bottleneck layer 40. A Multi-Layer Perceptron (MLP) is used in the supervised learning part to predict the hydrogen storage capacity. The major use of such an approach is to address the challenges of data scarcity and high feature dimensionality. In this study, the initial feature space has a high dimensionality and the available dataset is relatively small. As there are not too many machine learning studies available in the literature, it is not easy to choose proper descriptors. We have constructed 36 features using the Mendeleev software 41. To handle high-dimensional feature spaces with small datasets, we used an autoencoder to reduce the dimensionality of the data.
Figure 1: Architecture of our hybrid approach
2.1 Unsupervised learning
In the unsupervised learning part, we train an autoencoder on the entire dataset, which consists of only unlabeled data. The goal of the autoencoder is to learn a compact representation (or feature representation) of the data in the bottleneck layer.
Autoencoders are a kind of neural network that is used to extract features in an unsupervised manner. They are divided into two parts: an encoder, which compresses the input data into a lower-dimensional representation (also known as the latent representation or bottleneck), and a decoder, which reconstructs the original input from the compressed form 42.
The autoencoder’s purpose during training is to minimize the reconstruction error between the input and its reconstruction. This forces the encoder to learn a compact representation of the input data that maintains just the most significant information, which can then be utilized to extract features. Once trained, the encoder may be used to extract features from new, previously unseen data by feeding it through the network and using the bottleneck layer activations as features.
Let $\vec{x}$ be the input vector, and let $g_{\phi}$ and $f_{\theta}$ represent the encoder and decoder functions, respectively. The encoder function maps the input vector to a lower-dimensional representation $z$:
$z=g_{\phi}(\vec{x})=\sigma(W_{\phi}\vec{x}+b_{\phi})$ (1)
where $\sigma$ is the activation function, $W_{\phi}$ and $b_{\phi}$ are the weight and bias parameters of the encoder, respectively. For the activation function, we have used the RelU function as implemented in PyTorch 43.
The decoder function maps the lower-dimensional representation back to the original input space:
$\hat{x}=f_{\theta}(z)=\sigma(W_{\theta}z+b_{\theta})$ (2)
where $W_{\theta}$ and $b_{\theta}$ are the weight and bias parameters of the decoder, respectively.
The goal of the autoencoder is to minimize the reconstruction error, defined as the difference between the original input $x$ and its reconstruction $\hat{x}$:
$L(\phi,\theta)=||\vec{x}-\hat{x}||^{2}$ (3)
This loss function was minimized using the Adam optimization method to find the optimal values for the parameters $\phi$ and $\theta$ say $\phi^{*}$ and $\theta^{*}$ 44. Once the autoencoder is trained, the encoder function $z=g_{\phi^{*}}(x)$ can be used to extract features from the input data by using the activations of the lower-dimensional representation $z$ as the features.
The network architecture for both encoder and decoder includes two linear layers and dropout layers for preventing overfitting. The input to the autoencoder is first passed through a linear layer with 256 units, followed by a Leaky ReLU activation function 45. This is then followed by a dropout layer with a dropout probability of 0.3 to provide regularization.
As shown in Fig. 1, we create the elemental features using the Mendeleev software. These features are then fed into the autoencoder to learn latent information in an unsupervised manner. The importance of the latent space generation step lies in its ability to learn a more compact and relevant representation of the input features, potentially highlighting complex, nonlinear relationships between them that might be difficult to capture with a simple MLP. By training the MLP on these learned features instead of the original features helps to improve the prediction accuracy of the model. Also, as the dimension of the original feature is large it is very difficult to work with an MLP which is very which has five layers.
Given the unsupervised nature of the autoencoder training, it has the additional benefit of being less prone to overfitting, and more robust to noise and outliers in the input data, compared to directly training a supervised model on the raw input features.
2.2 Supervised learning
Let $z_{i}$ be the input data, and let $y_{i}$ be the corresponding target (true) value. $N$ is the total number of data points and $L$ is the number of labeled data points.
Here we train a regressor, such as a Multilayer Perceptron (MLP), on the labeled data using the extracted features through the above-mentioned method as inputs. The MLP takes the latent features as input and predicts the target values based on the labeled data. The Loss function to minimize can be represented as:
$L=\frac{1}{N}\sum_{i=1}^{L}(y_{i}-\hat{y}_{i})^{2}$ (4)
Here $\hat{y}_{i}$ is the prediction of the MLP, and N is the total number of data points.
2.3 ab-initio calculations
First-principles calculations were conducted employing the density functional theory (DFT) within the Vienna Ab initio Software Package (VASP) 46, 47. To determine the electronic energy, projected augmented wave (PAW) potentials and plane-wave basis sets were utilized for both core and valence electrons, employing the Perdew–Burke–Ernzerhof (PBE) functional 48, 49. The calculations accounted for long-range interactions using the semi-empirical Grimme D2 dispersion method and incorporated non-spherical contributions to the PAW potentials within the code 50. A cutoff energy of 500 eV ensured convergence of all energies. Structural optimization employed the conjugate gradient algorithm 51, with convergence criteria set at 10-7 eV for energy and 0.01 eVÅ-1 for force. Spin polarization was enabled in all calculations, except for the isolated H2 closed shell molecule.
3 Dataset
For this study, we utilized the Hydrogen Storage Materials Database shared by the HyMARC Data Hub (https://datahub.hymarc.org/). This is an openly available database that can be accessed at https://datahub.hymarc.org/en/dataset/hydrogen-storage-materials-db. This database contains information on more than 2000 hydride materials and their properties. The Hydride Database classifies the hydrides into eight classes: A2B intermetallic compounds, AB intermetallic compounds, AB2 intermetallic compounds, AB5 intermetallic compounds, complex hydrides, Mg alloys, solid solution alloys, and Misc. In the dataset, more than 950 compounds have 1-2 wthydrogen storage capacity (Fig. 2). The average hydrogen storage capacity of all compounds is 2.1 wt%. The dataset has as high as 20.8 wthydrogen storage capacity complex hydrides, which is BeB4H8 52.
Figure 2: Histogram showing the hydrogen weight percentage distribution in the database used for the study.
4 Description of the features
For every chemical composition, our feature generation algorithm creates a 36-dimensional vector that contains a set of nine elemental properties (Fig. 3). These characteristics include the atomic number, period, electronegativity (measured according to Pauling’s scale), electron affinity, atomic volume, atomic weight, fusion heat, and ionization energy. The covalent radius is calculated using Bragg’s method. Four statistical measures—the weighted sum, standard deviation, minimum, and maximum values—are computed within the compound’s composition for each property, resulting in $9\times 4=36$ dimensions.
Figure 3: Architecture of feature generation algorithm.
5 Results and Discussion
5.1 Machine Learning Model for the Prediction of Hydrogen Storage Capacity
In this study, our objective was to devise an effective machine learning model for predicting hydrogen storage capacity. Our investigation was bifurcated into two distinct scenarios: firstly, employing a Multilayer Perceptron (MLP) trained directly on all the features, and secondly, utilizing a hybrid approach.
5.2 Case I: Using the MLP Directly for All the Features
For our primary experiment, the MLP was trained and tested using all of the 36 features derived from the Mendeleev software. The dataset was strategically partitioned into training and testing subsets, with respective ratios of 0.8 and 0.2. The resultant outcomes of this methodology are visually represented in Fig.4. The predictive efficacy, quantified by the $R^{2}$ coefficient, was deemed unsatisfactory 53. Such results emphasize the inherent difficulties connected with managing datasets with a high number of dimensions, especially when the amount of data is small.
Figure 4: Pair plot contrasting the true values (from the dataset) with the predicted outcomes (both test and train) when the raw (36) features set is employed to train and test the MLP.
5.3 Case II: Adopting the hybrid Approach
Given the large number of dimensions present in the initial feature space (36 features) in contrast to the constrained size of the dataset, we opted to employ a hybrid approach. Our primary approach involved implementing an autoencoder to produce a latent space, which subsequently served as the input for a five-layer neural network (MLP). The objective of this downstream MLP was to predict the hydrogen storage capacity.
The optimal dimensions for the latent space were determined by assessing the model’s performance across various latent space sizes. The dataset was partitioned into training and testing datasets, adopting a 0.8 and 0.2 split ratio. The training process utilized a learning rate of 0.001 and spanned 1000 epochs.
Performance evaluation was undertaken by computing the $R^{2}$ value across different latent space dimensions. Our observations revealed that the $R^{2}$ value exhibited an initial increase with the expansion of the latent space dimension, post which it fluctuated (Fig. 5) The peak $R^{2}$ value (0.85) was obtained with a latent space dimension of $z=8$ (Fig. 6).
Figure 5: $R^{2}$ plotted against various latent space dimensions. A latent space dimension of $z=8$ yielded the most optimal fit for the MLP to the target.
Figure 6: Pair plot comparing the true values against the predicted values (both test and train) for a latent space dimension of $z=8$ and $R^{2}=0.85$
This outcome suggests that a latent space dimension of 8 effectively encapsulates the vast majority of information inherent in the original feature space, facilitating accurate predictions for hydrogen storage capacity. Thus, for ($z<8$) the latent space might be too small to capture all the relevant features of the input data. This can lead to underfitting, where the model fails to learn the underlying patterns effectively. The low dimensionality may force the autoencoder to compress the data so much that significant information is lost, making it difficult for the MLP to make accurate predictions. While for ($z>8$) the latent space might be too large, potentially leading to overfitting, where the model learns noise and details from the training data that do not generalize well to unseen data. A higher-dimensional latent space can represent the data more accurately, but it can also capture unnecessary or redundant information, which can degrade the performance of the MLP when it tries to generalize from the training data to new data.
6 Comparison with few other supervised models
In this section, we compare the performance of two supervised models, Lasso regression and Gradient Boosting regression, with the previously mentioned hybrid approach using an Autoencoder combined with a Multilayer Perceptron (MLP) 54, 55. We evaluate the models based on their R2 scores, which measure the proportion of the variance in the target variable that the models can explain. The results are shown in Fig.7. The R2 values obtained from the Lasso regression and Gradient Boosting regression are 0.48 and 0.8, respectively. These values indicate that both models have relatively lower predictive accuracy compared to the hybrid approach. The hybrid approach achieved an R2 score of 0.85, outperforming both linear models. These results suggest that the hybrid approach, combining an Autoencoder with an MLP, is more effective in capturing the underlying patterns and predicting the target variable compared to the individual linear models.
Figure 7: Performance of the Linear models (a) LASSO (b) GBR. Here the original 36 features are used.
6.1 Identification of new hydrogen storage materials
To find new materials for efficient hydrogen storage, we used the following approach: We calculated the Pearson correlation function for different physically meaningful features, including the target - the storage capacity (Wt), which we show in Fig. 8 56. We take the weighted sum for all nine features and compute the correlation. It can be seen that the period number and electronegativity have the highest correlation with the weight percentage. A significant Pearson correlation between the period number and electronegativity with the target hydrogen storage capacity, suggests their pivotal roles in influencing a material’s ability to store hydrogen efficiently. The period number, reflecting the electron shell structure, directly impacts its hydrogen bonding capabilities, while electronegativity, indicative of an electron attraction strength of the material, influences the stability and capacity of hydrogen storage in metal hydrides. Therefore, we first scanned the metal hydrides from the Materials Project with these two features. Next, we constructed a 36-dimensional feature space using the above materials, as before (using four statistical measures for each one). Then, we used pre-trained embeddings from the autoencoder so that the property (hydrogen weight in percentage) can be computed using the MLP. Finally, we performed DFT-based calculations to re-verify these values.
Figure 8: Correlation between different physically meaningful features and the target. Here atomic number ($Z$), period ($P$), electronegativity ($\chi$), electron affinity ($\chi_{a}$), atomic volume ($V_{a}$), atomic weight (W), fusion heat ($H_{f}$), and ionization energy ($IE$) is considered.
Using the above approach we identified two materials Al11O18 and V2O5. Both the materials are stable from our formation energy calculations. The DFT calculated formation energies are respectively -3.34 eV/atom and -2.12 eV/atom for Al11O18 and V2O5. From the DFT approach the hydrogen storage capacities were calculated using the hydrogen adsorption energies 11 shown below,
$E_{ads}=\frac{1}{n}\times\left(E_{structure+nH}-\left(E_{structure}+\frac{n}{2}E_{H_{2}}\right)\right)$ (5)
where Estructure+nH is the energy of the hydrogenated structure, Estructure and EH2 are the energies of the pristine structure and isolated H2, respectively, and n is the number of hydrogen atoms involved in the adsorption.
To predict hydrogen storage capacity, all interstitial sites of Al11O18 and V2O5 are gradually filled with hydrogen, and the adoption energy is calculated 57. Figure 9 shows the change in adsorption energy with increasing hydrogen concentration. The adsorption energy increases as the hydrogen fraction increases. The maximum hydrogen fraction at which adsorption energy remains negative is used to predict hydrogen storage capacity. The predicted hydrogen storage capacities from the DFT calculations are respectively 4.61% and 3.83% while from the MLP model, we obtain values respectively 5.98% and 4.35%.
Figure 9: Adsorption energy of hydrogen in Al11O18 (a) and V2O5 (b) at different H concentration.
Figure 10: The total (top panel) and partial DOS (down panel) for the predicted materials Al11O18 V2O5 and their hydrides. The Fermi level is set at zero energy and marked by the vertical lines.
In the Fig.10, we show the density of states (DOS) for the machine learning predicted materials and their hydrides for different hydrogen loadings. It was found that Al11O18 shows a more stable behavior when loaded with hydrogen compared to V2O5 because V2O5 relatively DOS at the Fermi energy for all the hydrogen loading. For both materials, H-s strongly hybridizes with the O-p states as can be seen from the projected DOS of Al11O18H4 and V2O5H shown in the down panel of the above figure. In the case of V2O5H, the anti-bonding states formed between V-d and H-s are also filled and therefore the system shows less stability.
6.2 Use of Partial Weight-Transfer and Freezing Encoder Weights
So far, we have discussed the prediction of hydrogen weight percentage by directly feeding the latent features ($z$) to the MLP without considering the trained weights of the autoencoder. One might be interested to know what if along with the latent features the pre-trained weights from the autoencoder are also transferred in the downstream MLP task. In the supervised learning phase, we explored a hybrid approach that integrates weights from the trained autoencoder into the MLP for predictive modeling.
The autoencoder follows the architecture of dim_in $\rightarrow$ 256 $\rightarrow$ dim_latent $\rightarrow$ 256 $\rightarrow$ dim_in, encoding and then decoding the information. dim_in is the original feature dimension (36), while dim_latent is the dimension of the latent space. In contrast, the original MLP architecture is dim_latent $\rightarrow$ 512 $\rightarrow$ 256 $\rightarrow$ 16 $\rightarrow$ 1, a straightforward feedforward network without the encoding-decoding mechanism.
To facilitate transfer learning and leverage the learned representations from the autoencoder, we adjusted the first two layers of the MLP to align with the autoencoder’s encoder structure. The modified MLP structure is dim_latent $\rightarrow$ 256 $\rightarrow$ dim_latent (8) $\rightarrow$ 512 $\rightarrow$ 256 $\rightarrow$ 16 $\rightarrow$ 1. This structure begins by expanding dim_latent to 256, then contracts back to dim_latent (8), before proceeding through the remaining MLP layers.
Figure 11: Schematic representation of partial weight transfer from the Autoencoder to the MLP. The Autoencoder encodes the input dimension into a latent space ($z$) and then decodes back to the original dimension. In the MLP, the first layer expands $z$ to 256, contracts back to $z$, and then proceeds through additional layers to reach the final single output (storage capacity). The red dashed line indicates the transfer of weights from the Autoencoder’s encoder to the initial layer of the MLP.
As depicted in Fig.11, the process of weight transfer involves the transfer of weights from the encoder part of the autoencoder, specifically the weights leading into the latent space $z$, which are copied to the corresponding layer of MLP.
In addition, in order to safeguard the capability of the autoencoder to extract features, we maintained the weights of the transferred encoder layers in a frozen state throughout the training of the MLP. This was accomplished by ensuring that their weights remain unaltered during the backpropagation phase. By adopting this approach, the MLP is able to benefit from the pre-trained representations in the autoencoder while also acquiring the ability to learn features specific to the given task in subsequent layers. The amalgamation of weight transfer and selective weight freezing constitutes an innovative method in our modeling process. However, despite the presence of some improvements, it cannot be deemed as remarkable. Keeping the remaining parameters unchanged, we obtained an R-squared value of 0.86 for z=8, which represents a modest improvement when compared to our previous scenario where only the latent features were transferred to the MLP. Thus this study highlights the complexities and limitations of applying transfer learning techniques in the context of materials science. One possibility in this particular case off-course could be that the training of the MLP with fixed encoder weights may give rise to optimization difficulties, wherein the propagation of gradient updates across the network may not be efficient due to the unchangeable nature of specific weights. The other possibility could be that the initial task for which the autoencoder was trained may not align well with the new task in terms of weight, meaning that the weights learned may not be optimal for predicting the hydrogen weight percentage even though features were useful.
7 Application to the Hydrogen Storage Materials Generated Using a Large Language Model
To show further the predictive capability of our approach, we apply it further to a set of materials that we generate using a large language model (LLM). Leveraging the power of GPT-2 (Generative Pre-trained Transformer version 2), a model renowned for its proficiency in natural language processing, we have generated a few new hydrogen storage materials. Unlike the resource-intensive traditional methods such as Generative Adversarial Networks (GANs) or diffusion models, which have limited testing in materials science, our approach utilizes a fine-tuned GPT-2 model, offering a more quick and targeted pathway to innovation. The details of the methodology are explained as detailed by Fu et. al 58 which leverages the advanced capabilities of LLMs originally developed for natural language processing, adapting them to the intricate language of material compositions. The fine-tuning process involves adapting a pre-trained GPT-2 model to the specific nuances of material science. This adaptation is facilitated by a custom vocabulary. which includes a comprehensive list of chemical elements and special tokens. This specialized vocabulary is crucial for accurately representing and processing material compositions. Our training process fine-tunes the model with a dataset specifically curated for hydrogen storage materials, allowing the model to learn and predict the complex patterns of material composition. The cornerstone of this methodology is the training of these models on comprehensive databases of known material compositions, including the ICSD (Inorganic Crystal Structure Database), OQMD (Open Quantum Materials Database), and Materials Project. This extensive assimilation of existing material data equips LLMs like GPT-2 with the ability to understand and replicate the intricate patterns and rules that govern material compositions. Through this process, the models are enabled to generate new, chemically plausible material compositions, which hold great potential for hydrogen storage applications. In an extension to the work by Fu et al. 58 on materials generation, we take a slightly modified approach where not only the chemical composition of materials but also a specific property: their hydrogen storage capacity is also considered to generate the chemical labels.
Figure 12: Flowchart depicting the training and generation processes for a GPT-2 language model. The left column (Training Phase) outlines the steps from initializing parameters to saving the trained model. The right column (Generation Phase) demonstrates the sequence generation process using the trained model.
As a result, the training of the model emphasizes the detailed relationship between a material’s composition and its storage capacity. Such integration enhances the model’s capacity to predict materials more accurately. The configuration of the model is tailored for this task, with an embedding size of 180, 2 attention heads, and 6 layers, which makes for a robust yet effective learning structure. Furthermore, the model’s vocabulary has been expanded to include numerical values, facilitating the inclusion of hydrogen storage capacity data.
Starting from a sequence of input tokens (comprising chemical symbols and storage capacity) the model first computes their embeddings. It then applies learned weight matrices to derive the usual queries, keys, and values matrices $Q$, $K$, and $V$, which are used to calculate the self-attention scores given by $\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ 59 where $d_{k}$ is the dimensionality of the key vectors. These scores determine how much each token in the sequence should attend to every other token, capturing both local and global dependencies within the sequence. During the training phase, the objective is to fine-tune a model’s parameters, denoted by $\theta$, to accurately predict material compositions from a given sequence of elements. This process involves presenting the model with sequences of tokens $[t_{1},t_{2},\ldots,t_{n}]$, each representing either a unique element within a material composition or a numerical value representing the storage capacity. The model learns by adjusting $\theta$ to maximize the log-likelihood of correctly predicting the next token $t_{n+1}$ in the sequence, formalized as $\max_{\theta}\sum\log P(t_{n+1}|t_{1},\ldots,t_{n};\theta)$. Here, $P(t_{n+1}|t_{1},\ldots,t_{n})$ represents the conditional probability of observing the next token $t_{n+1}$ given the preceding sequence $[t_{1},\ldots,t_{n}]$. The self-attention mechanism within the model facilitates this learning process by allowing each token to dynamically influence the prediction of the next token based on its context and relevance to the rest of the sequence.
During generation, the trained model is loaded to predict new material compositions. Starting with a random seed $S_{\text{rand}}$, which may consist of one or more element tokens and a random value of the storage capacity, the model iteratively generates the next element in the sequence by sampling from the probability distribution $P(t_{n+1}|t_{1},\ldots,t_{n};\theta)$ it learned during training. This process continues, adding each newly predicted element to the growing composition, until a termination condition is met, such as reaching a predefined sequence length or generating a special end-of-sequence token. The generation process is informed by the patterns and rules learned during training (or precisely the grammar of the chemistry), enabling the exploration of new and potentially viable material compositions that were not present in the original training dataset.
Using this methodology, we have successfully generated 225 new materials with potential applications in hydrogen storage.
Materials No of H FE (eV) $\mu_{H}^{C}$ $\mu_{H}^{X}$ Wt(DFT) Wt(ML) Eh Expt.
$)H2C2N$+6 -0.11 -3.5 -3.3 10.06 10.8 0.46 yes
[…]
$\mathrm{MgC}$ 2 0.37 -3.0 -3.3 5.25 4.65 1.25 no
$\mathrm{NLi}$ 3 -1.67 -5.05 -3.3 6.73 4.5 1.34 yes
Table 1: Examples of hydrogen storage materials identified by GPT-2 model. From the Materials project, we report the structures with the lowest energy above Hull Eh (eV), whether they are experimentally observed or not. From the hybrid method mentioned above (Wt(ML)), we calculated their storage capacity as well and compared it with the corresponding DFT-values (Wt(DFT)). The last column indicates whether the material is experimentally synthesized (yes) or not (no).
Here the critical hydrogen chemical potential, $\mu_{H}^{C}$ is defined as $\mu_{H}^{C}=1/n\left(E_{MAX+nH}-E_{MAX}\right)$, which is lower than the chemical potential $\mu_{H}$ of the hydrogen at gas phase at very low temperature (as our calculation corresponds to DFT temperature scale). ”FE” stands for the formation energy. From the above table, it is apparent that our model exhibits commendable performance and is quite reliable for the calculation of the hydrogen storage capacity. In the final table 2, we present an additional five materials that have been predicted by GPT-2 and whose hydrogen weight percentage has been computed by our model. As the atomic configurations of these materials remain elusive, we employ the methodology expounded by Kusaba et al 60, which employs the approach of metric learning, where the algorithm has been trained on a substantial amount of crystal structures that have already been identified to ascertain the isomorphism of crystal structures formed by two specified chemical compositions, yielding an accuracy of approximately 96.4%. For a given composition with an unknown crystal structure, this approach automatically selects a collection of template crystals from a database of crystal structures that possess nearly identical stable structures. When we subject the compositions predicted by GPT-2 to this method, we obtain a collection of structures for each composition. We consider the one with the highest spacegroup and then perform the structural optimization followed by the calculation of the storage capacity using our hybrid-ML approach. The rightmost column of the table visually represents the crystal structure of each material.
Material Wt(ML) Spacegroup Structure
NCAl 7.15 164 [Uncaptioned image]
MgCHF 6.75 186 [Uncaptioned image]
CAlB 6.24 194 [Uncaptioned image]
MgCF 5.69 129 [Uncaptioned image]
B2NLi 7.18 166 [Uncaptioned image]
Table 2: Materials predicted by GPT-2 model and their storage capacities obtained using our hybrid ML approach. The spacegroup information was obtained via DFT optimization of the structure with the highest spacegroup predicted by the approach mentioned in Kusaba et al 60
It will be worth synthesizing experimentally these materials for enhanced hydrogen storage applications. In future work, we plan to explore the use of advanced models like GPT-3.5 which could be highly beneficial. GPT-3.5, with its sophisticated natural language understanding and generation capabilities, can potentially enhance the prediction and design of novel hydrogen storage materials. Its advanced algorithms and large knowledge base might lead to more accurate predictions and innovative approaches in material science, opening new directions for efficient and effective hydrogen storage solutions.
8 Conclusion
This study investigated the efficacy of a hybrid deep learning framework for predicting hydrogen storage capacities in metal hydrides, addressing the critical issue of limited experimental data. By leveraging an advanced autoencoder and elemental descriptors from Mendeleev software, we effectively tackled the challenge of high-dimensional feature spaces inherent to materials informatics, where data is often sparse and heterogeneously distributed. A deep Multi-Layer Perceptron (MLP) architecture with five layers served as the prediction model, utilizing latent representations generated by the autoencoder to capture crucial features from high-dimensional data. The framework was tested for a few materials that were selected as potential hydrogen storage materials from the features that are highly correlated with the target as well as for the materials that are generated by a large language model: GPT-2. In both cases, the predicted capacity by the ML approach matches quite well with the DFT calculations. We further discuss the issue of weight transfer from the Autoencoder to the MLP in addition to the latent features.
Acknowledgements
This work was supported by the Korea Institute of Science and Technology (Grant number 2E31851), GKP (Global Knowledge Platform, Grant number 2V6760) project of the Ministry of Science, ICT and Future Planning.
9 Conflict of interest
The authors declare no conflict of interest.